OpenStudy (hotpinkgirl):

during exercise the recommended maximum heart rate in beats per minute is modeled by the formula m = 176 - 0.8A, where M is the maximum heart rate and A is the persons age. aproximately, at what age would you have recommended heart rate in beats per minute of 135? A. 25 years B. 39 years C. 51 years D. 53 years

OpenStudy (hotpinkgirl):

@texaschic101 @SmoothMath @EclipsedStar @freckles @jagr2713 @KyanTheDoodle @sasogeek @dan815

OpenStudy (korosh23):

Ok now we know that m is the rate right?

OpenStudy (korosh23):

@HotPinkGirl

OpenStudy (hotpinkgirl):

right

OpenStudy (korosh23):

Ok so in the question it says m is 135. Now we know that the rate is 135

OpenStudy (korosh23):

We plug in the number

OpenStudy (korosh23):

\[135= 176 - 0.8A\]

OpenStudy (korosh23):

This is the equation we want to solve for A which is the age of the person.

OpenStudy (korosh23):

Are you understanding until now? :)

OpenStudy (hotpinkgirl):

no

OpenStudy (korosh23):

What is the thing that is confusing you

OpenStudy (hotpinkgirl):

how do we get the answer?

OpenStudy (korosh23):

Oh but I still did not finish yet. Did you understand that m is the rate so we have to replace it with 135?

OpenStudy (hotpinkgirl):

yes

OpenStudy (korosh23):

Ok awsome let's move on

OpenStudy (korosh23):

\[135= 176 - 0.8A\] \[135 - 176 = 176 - 0.8A - 176\]

OpenStudy (korosh23):

Ok now we want to only get A, so we have to take 176 to the other side, which becomes negative. ALWAYS remember that when you do something to one side you have to do it to the other side as well. For example: x +2 = -4 x+ 2 -2= -4 -2 x= -6

OpenStudy (korosh23):

Do you understand?

OpenStudy (korosh23):

@HotPinkGirl

OpenStudy (hotpinkgirl):

yes

OpenStudy (korosh23):

Awsome! :D

OpenStudy (korosh23):

ok after doing \[135 - 176 = - 0.8A\] Now we do the left side which we get \[-41= -0.8A\]

OpenStudy (korosh23):

ok do you understand this as well?

OpenStudy (hotpinkgirl):

yes

OpenStudy (korosh23):

\[-41= -0.8A\]

OpenStudy (korosh23):

What should we do now? Remember when we do something to oneside, we have to do it to the other side as well.

OpenStudy (korosh23):

@HotPinkGirl are you with me?

OpenStudy (hotpinkgirl):

yes

OpenStudy (korosh23):

so what should we do know?

OpenStudy (hotpinkgirl):

now we do the other side

OpenStudy (korosh23):

ok how?

OpenStudy (korosh23):

\[-41= -0.8A\] We are solving for A. That means A has to stay and does not change. Thus, what number changes?

OpenStudy (hotpinkgirl):

-41

OpenStudy (korosh23):

-41 changes, but for the right side we have -0.8A. What number should be removed?

OpenStudy (hotpinkgirl):

0.8

OpenStudy (korosh23):

-0.8 well done. Do not forget the negative sign. Now we have to do it like this. \[\frac{ -41 }{ -0.8 }= \frac{ -0.8A }{ -0.8 }\]

OpenStudy (korosh23):

The rule is when we do something to one side we have to do it to the other side as well.

OpenStudy (korosh23):

Now A= ?

OpenStudy (hotpinkgirl):

51 years?

OpenStudy (korosh23):

You are right. Perfext. :D @HotPinkGirl if you liked my answer please give me a medal. I would appreciate if you write me a fan testimonial as well. Thank you!

OpenStudy (hotpinkgirl):

ok, thanks

OpenStudy (korosh23):

Your welcome. :)